Solve the following systems of equations:

2x - 2y - 2z = -2
-x + y + 3z = 0
-3x + 3y - 2z = 1

Answer by stanbon(75887)

Solve the following systems of equations:
2x - 2y - 2z = -2
-x + y + 3z = 0
-3x + 3y - 2z = 1
-------------------------
Modify:
x - y - z = -1
-x + y + 3z = 0
-3x + 3y - 2z = 1
-------------------------
Add 1st to 2nd ; Add 3 times 1st to 3rd to get:
x - y - z = -1
0 + 0 + 2z = -1
0 + 0 - 5z = -2
--------------------
z = -1/2 and z = 2/5
This is a contradiction.
There is no system solution; at least two of the planes are parallel.
=======================================================================
Cheers,
Stan H.